OpenStudy (anonymous):

y - 5x = -6 and 3y - 15x = -12

OpenStudy (anonymous):

a. one b. two c. infinitely many d. none

OpenStudy (anonymous):

@courter98

OpenStudy (anonymous):

How do you think you start?

OpenStudy (anonymous):

-6 + 5x = y

OpenStudy (anonymous):

wait I think I might have it, hold on

OpenStudy (anonymous):

You have to get the two equations to be divisible by the same number. Since in the second equation, they are all divisible by 3. You would multiply the first equation by 3

OpenStudy (anonymous):

y - 15

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

6 + 5 x 3 = 15 I honestly don't know how to do this, im trying to study for semester and unit exams. can you show step by step on how to do it

OpenStudy (anonymous):

Sure. Give me one sec

OpenStudy (anonymous):

y - 5x = -6 and 3y - 15x = -12 have to both be divisible by the same number. 3y - 15x = -12 are all divisible by 3.

OpenStudy (anonymous):

So to get y - 5x = -6 divisible by 3 you have to multiply all of them by 3. y(3) -5x(3) = -6(3)

OpenStudy (anonymous):

Do you understand so far?

OpenStudy (anonymous):

yes, i'm writing this down

OpenStudy (anonymous):

Sorry my internet cut out

OpenStudy (anonymous):

You should have gotten 3y - 15x = -18

OpenStudy (anonymous):

my computer is acting weird, you can continue

OpenStudy (anonymous):

3y - 15x = -18 3y - 15x = -12 Now all you do is subtract

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The y's and the x's cancel out. so you are left with 0 = -6. 0 = -6 is never true. Therefore, the answer is no solutions

OpenStudy (anonymous):

thanks so much, I hope I do well with these type of questions on the exams

OpenStudy (anonymous):

I wish you the best of luck! Is this algebra 1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok. If you need any more help, just let me know. I did really good in algebra 1